# Semantics-Empowered Space-Air-Ground-Sea Integrated Network: New Paradigm, Frameworks, and Challenges

### v1

@@ I-C Contributions
To fill the void in literature of semantic communications for the SAGSIN, in this survey, we propose a comprehensive paradigm called PCCAIP which fully addresses fundamental limitations in the SAGSIN. Sequentially, we review recent literature on semantic communications from three perspectives, which are 1) significance representation and protection, 2) meaning similarity measurement and meaning enhancement, and 3) ultimate effectiveness and effectiveness yielding. Specifically, our contributions are listed as follows.
@@
We propose a new paradigm for the implementation of semantic communications in the SAGSIN called PCCAIP, in which perception, computing, and actuation techniques will be jointly utilized for communications, and thus the design for the three types of state-of-the-art semantic communication systems can be guided by PCCAIP.
@@
I-D Organization
The rest of this survey is organized as follows. Section II propose the PCCAIP, where utilization of perception, computing, and actuation modules in the SAGSIN is elaborated respectively. In Section III, we introduce significance metrics as semantic measures, and review SPTs including significance-oriented sampling, coding, and modulation. Section IV introduce meaning similarity metrics, DL models, and METs based on meaning extraction and reconstruction. The focus of Section V is task-oriented communications guided by PCCAIP, where effectiveness metrics are reviewed and EYTs are discussed for specific tasks. We then envision some future challenges on semantic communication implementation in the SAGSIN in Section VI, and VII concludes the survey. The architecture of this survey is illustrated in Fig. 2.
@@ II Perception, Computing, and Actuation for Semantic communications in the SAGSIN
As mentioned in Section I, semantic communication systems play a crucial role in the SAGSIN by enabling efficient information exchange and intelligent actuation. However, relying solely on traditional communication techniques may not meet the requirements of complex tasks due to the limitations in network resources. Therefore, perception, computing and actuation techniques facilitate semantic communications to form the PCCAIP. In the following discussion, we will firstly address these techniques, and then provide a detailed introduction to the PCCAIP.
@@ II-A Perception
The integration of space-based, air-based, ground-based, and sea-based networks results in a significant increase in business volume. Multiple types of sensors deployed in these networks enable real-time perception of targets while generating a substantial amount of multimodal data, which pose a great challenge in supporting services and applications with extremely low latency requirements. One feasible solution is to reduce the volume of data to decrease the latency in subsequent communication and processing stages. This is why we need to perceive the significance, meaning, and effectiveness-related information of the source data in the SAGSIN. Capturing significance can be achieved through significance-oriented sampling, while capturing the other two can be accomplished through intelligent semantic extraction, which is typically facilitated by DL, knowledge graphs (KGs), and so on. The significance-oriented sampling and semantic extraction are collectively referred to as perception in this survey. Perception plays a significant role in various application scenarios. For example, in the emergency response and rescue applications for maritime incidents, ocean buoys deployed in the maritime area are used to collect oceanographic source data such as sea currents, wind speeds, and wave conditions. UAVs deployed above the sea surface capture image data with cameras. During non-urgent situations, the sensors operate at a relatively low sampling frequency to conserve energy and reduce data processing burden. However, in the emergency events such as maritime shipwrecks, oil spills, fires, or drowning, the sensors immediately increase their sampling frequency to keep data freshness and let the server estimate the source status more accurately. The sampled data are further processed to extract semantic information, such as the situation of stranded crew members and ocean current information, which is then transmitted to the emergency rescue center in real-time to facilitate rescue operations.
@@ II-B Computing
In the semantic-empowered SAGSIN, computing refers to the AI-related complex operations, such as the feature extraction, semantic coding, and adaptive decision. Therefore, it is imperative to use AI algorithms to extract semantic information and make intelligent decisions, which will release data traffic and thus decrease processing latency. However, AI algorithms rely heavily on sufficient computing resources, and it is unlikely to achieve satisfactory performance when these algorithms are deployed on end devices such as ocean buoys and UAVs due to limitations in computing and caching resources. The SAGSIN encompasses various types of devices, resulting in a large volume of data that usually requires real-time processing and analysis. Edge computing technology has emerged to address this issue. In recent years, there have been significant academic and industrial interests in on-orbit edge computing based on space-based networks[29, 30], as well as edge storage technologies [31, 32]. A large number of cloud servers, edge servers, and other computing units are distributed in the SAGSIN, providing supports for perception, communication, and actuation in various stages. For instance, in the context of satellite-based edge computing, efficient interconnection is achieved among space-based, and ground-based networks by leveraging LEO satellites as the core. An edge computing platform is built on the satellites, empowering the entire network with on-orbit capabilities for intelligent data collection, processing, and caching[33]. The authors assume a geographically dispersed remote team with diverse members. Through augmented reality (AR) and virtual reality (VR) technologies, remote team members can share a virtual workspace, enabling collaborative working and real-time communications. In this scenario, the on-orbit edge servers are equipped with a semantic encoding network driven by DL models. Team members wear head-mounted displays (HMD) integrated with semantic decoding networks, accessing the shared virtual workspace through VR technology. In this example, the satellite-based edge computing platform provides robust support for the combination of VR technology and semantic communications.
@@ II-C Actuation
The receiver utilizes the received messages, which are relevant to task performance, to accomplish the task decision. Based on the decision results, the task execution process will exert an influence on the physical world. In this survey, the task decision and task execution comprise the definition of actuation. Taking remote control services as example, the intelligent decision-maker generates and transmits task decision commands based on the received messages, and then the terminal receives the decision commands and executes specific tasks. With the development of 5G/6G, the variety of task-critical applications in the SAGSIN continues to enrich. For instance, the stability control of UAVs can be modeled as a classical CartPole problem solution. In this scenario, sensors collects environmental data, which will be then processed on the edge server using DL-based autoencoders to extract task-related semantic features. These features are transmitted to the remote control center on the ground to generate control commands, enabling remote control of the UAVs [34]. Another example is found in remote surgeries, which is a typical use case of haptic Internet. Sensors located in remote areas (e.g., mountainous areas and remote seas) collect crucial patient data, which are transmitted to the medical center with the aid of satellites. Surgeons receive real-time auditory, visual, and haptic feedback through a human-machine interface (HMI). Visual feedback is provided using streaming technologies, such as holographic-type communications (HTC), depending on whether surgeons interact with holograms or wear head-mounted devices. Subsequently, surgeons operate haptic devices through the HMI and perform surgical actions based on real-time visual feedback and haptic information transmitted to the robot[35].
@@ II-D PCCAIP
According to the interpretation of perception, computation, actuation, we introduce the perception-communication-computing-actuation-integrated framework under the guidance of PCCAIP. This framework aims to maximize the effectiveness of tasks by establishing closed information flow loop, as shown in Fig. 3. Specifically, we consider an emergency detection and rescue scenario in the SAGSIN as an example. Various sensors (e.g., cameras deployed on the UAVs) continuously observe the environment that contains abnormal statuses (e.g., sinking ships or drowning persons), and decide to sample data representing the current statuses when the data are significant enough (e.g. untimeliness, status change, or environment change). After data sampling, these data will be further processed by the computing-intensive semantic extractor deployed on UAVs and then transmitted to remote edge servers (e.g. servers deployed on maritime base station or space base station), by which multi-level data perceptions are completed by distributed computing resources. Based on the received status data, the edge servers conduct real-time decision of whether or not the terminal (e.g., rescue ship) rescue the emergencies and which emergencies are to be rescued (if rescue abilities are limited), and then transmit the decision commands to remote terrestrial cloud server via LEO/GEO satellite relays. After the cloud server received the decision commands, it instantly conduct task execution (e.g., rescuing the ships/persons, or doing nothing) based on the commands, by which task actuation is completed. Task execution will further determine the statuses and thus the decisions at the next time interval (for example, if a sinking ship gets rescued by the rescue ship, there will be no emergency in the next period of time, and thus the rescue ship does not need to rescue in the next), by which closed-loops for the emergency rescue tasks are formed by the PCCAIP framework.
@@ II-E Lessons Learned from This Section
This section elaborates on how the perception, computing, and actuation techniques help realize semantic communications. Overall, perception techniques help select the data with more semantics, computing techniques facilitate intelligent communications in various aspects, and actuation techniques aid in the interaction between system and environment to complete all-inclusive tasks. PCCAIP integrates all these techniques which can realize a closed-loop communication and a joint design to achieve high effectiveness of tasks. Moreover, maritime rescue tasks are adopted as an example to showcase how PCCAIP guides semantic communication system design. As typical semantic communication frameworks facilitated by PCCAIP, we will introduce significance-oriented, meaning-oriented, and effectiveness/task-oriented communications systems respectively in the following sections. In Section III, we design a significance-oriented semantic communication system which belongs to a perception-communication-integrated semantic communications. In Section IV, we study a meaning-oriented semantic communication system which falls under the category of a perception-communication-computing-integrated semantic communication system. In Section V, we introduce a task-oriented communication system which achieves perception-communication-computing-actuation integration.
@@
In this section, we will review the research on semantic communications from the data significance perspective. Here, the term “semantics” is explained as its etymological meaning “significance”, which implies a significance-oriented semantic communication system. In significance-oriented semantic communications in the SAGSIN, a joint perception-communication framework is adopted as shown in Fig. 4. Specifically, source data are firstly sampled by significance-oriented sampler in order to release the heave data burden, and the sampled data will be coded and modulated for transmission over wireless channel. The receiver demodulates and decodes the received message to recover sampled data. According to the recovery results, the significance of source data varies to affect the evolution of significance metric; sequentially, the variation of significance metric will affect the further decision of sampler and coder/modulator at the transmitter via feedback from the receiver. It is worth noting that both coding and modulation techniques mentioned in this section are traditional technologies based on Shannon information theory. Considering the fundamental limits of the SAGSIN and the time-sensitivity of various tasks facilitated by SAGSIN, the data freshness/timeliness is naturally significant for semantic communication system due to large communication distance. Therefore, timeliness metrics, including AoI and its variants, are of vital importance in describing the performance of the tasks. Thus, we will first introduce AoI and AoI variants as significance metrics in subsection III-A, and then review the literature on techniques towards significance-protection in subsection III-B.

### v3

@@ I-D Contributions
To fill the void in literature of semantic communications for the SAGSIN, in this survey, we propose a comprehensive paradigm called PCCAIP which fully addresses fundamental limitations in the SAGSIN. Sequentially, we review recent literature on semantic communications from three perspectives, which are 1) significance representation and protection, 2) meaning similarity measurement and meaning enhancement, and 3) ultimate effectiveness and effectiveness yielding. The contributions of the related survey papers and ours are compared in Table I. Specifically, our contributions are listed as follows.
@@
We propose a new paradigm for the implementation of semantic communications in the SAGSIN called PCCAIP, in which perception, computing, and actuation techniques are jointly utilized for communications, and thus the design for the three types of state-of-the-art semantic communication systems can be guided by PCCAIP.
@@
I-E Organization
The rest of this survey is organized as follows. Section II proposes the PCCAIP, where utilization of perception, computing, and actuation modules in the SAGSIN is elaborated respectively. In Section III, we introduce significance metrics as semantic measures, and review SPTs including significance-oriented sampling, coding, and modulation. Section IV introduces meaning similarity metrics, DL models, and METs based on meaning extraction and reconstruction. The focus of Section V is task-oriented communications guided by PCCAIP, where effectiveness metrics are reviewed and EYTs are discussed for specific tasks. We then envision some future challenges on semantic communication implementation in the SAGSIN in Section VI, and VII concludes the survey. The architecture of this survey is illustrated in Fig. 2, and the majority of the abbreviations/acronyms used in this survey are listed in Table II.
Abbreviation | Full Name | Abbreviation | Full Name | ||
5G/6G | fifth/sixth generation | SAG(S)IN |
LEO/GEO | low/geostationary Earth orbit | UAV | unmanned aerial vehicle | ||
HAP | high-altitude platform | IoT/IoV | Internet of things/vehicles | ||
LDPC | low density parity check | NOMA | non-orthogonal multiple access | ||
AI | artificial intelligence | DL | deep learning | ||
MSE/MAE | mean square/absolute error | JSCC | joint source-channel coding | ||
SPT | significance protection technique | MET | meaning-enhancement technique | ||
EYT | effectiveness-yielding technique | AR/VR/MR | augmented/virtual/mixed reality | ||
AoI | age of information | AoS | age of synchronization | ||
AoII | age of incorrect information | UoI | urgency of information | ||
GoT | goal-oriented tensor | QAoI | query age of information | ||
VoI | value of information | ACK | acknowledgment | ||
MDP | Markov decision process | (H)ARQ | (hybrid) automatic repeat request | ||
HARQ-CC | HARQ with chase combining | (HARQ-)IR | (HARQ with) incremental redundancy | ||
BSC/BEC | binary symmetric/erasure channel | SNR | signal-to-noise ratio | ||
AWGN | additive white Gaussian noise | FBL | finite block-length | ||
NACK | negative acknowledgment | OMA | orthogonal multiple access | ||
MDP | Markov decision process | PSO | particle swarm optimization | ||
ANN | artificial neural network | DNN | deep neural network | ||
NN | neural network | BER | bit error rate | ||
WER | word error rate | BLEU | bilingual evaluation understudy | ||
CIDEr |
MSS | metric of semantic similarity | ||
BERT |
PSNR | peak signal-to-noise ratio | ||
SSIM | structural similarity index | MS-SSIM | multi-scale SSIM | ||
LPIPS |
GAN | generative adversarial network | ||
FID/FDSD | Fréchet Inception/DeepSpeech distance | KID/KDSD | kernel Inception/DeepSpeech distance | ||
AKD | average keypoint distance | SDR | signal-to-distortion ratio | ||
PESQ | perceptual evaluation of speech quality | IRS | intelligent reflecting surface | ||
FFT | fast Fourier transform | MCD | Mel cepstral distortion | ||
RL | reinforcement learning | SSCC | separate source-channel coding | ||
DeepJSCC | deep-learning based JSCC | CNN | convolutional neural network | ||
RNN | recurrent neural network | ReLU | rectified linear unit | ||
FCN | fully convolutional network | LSTM | long short-term memory | ||
GRU | gated recurrent unit | NLP | natural language processing | ||
GPU | graphics processing unit | GCN | graph convolutional network | ||
RLN | reinforcement learning network | DeepSSCC | DL-based SSCC | ||
CSI | channel state information | BPSK | binary phase shift keying | ||
NTSCC | nonlinear transform source-channel coding | RS (codes) | Reed-Solomon (codes) | ||
KG | knowledge graph | KB | knowledge base | ||
ProbLog | probabilistic logic programming language | ML | machine learning | ||
TP/TN | true positive/negative | FP/FN | false positive/negative | ||
IoU | intersection over union | WNCS | wireless network control system | ||
AoTI | age of task information | SNN | spiking neural network | ||
(Dec-)POMDP | (decentralized) partially observable MDP | VQA | visual question answering |
@@ II Perception, Computing, and Actuation for Semantic communications in the SAGSIN
As mentioned in Section I, semantic communication systems play a crucial role in the SAGSIN by enabling efficient information exchange and intelligent actuation. However, relying solely on traditional communication techniques may not meet the requirements of complex tasks due to the limitations in network resources. Therefore, perception, computing and actuation techniques facilitate semantic communications to form the PCCAIP, by which a ”target-perception-communication-actuation-target” closed-loop framework facilitated by intelligent computing is constructed. Specifically, perception techniques are utilized to observe the targets (e.g., abnormal statuses shown in Fig. 3) in the physical world. Next, communication techniques help transmit the observed target information to remote nodes (e.g., edge servers in Fig. 3) for further processing. Sequentially, actuation techniques will guide the actuator (e.g., both edge servers and cloud servers in Fig. 3) to generate the best task decision commands. Finally, the commands will be executed by the terminal, and the status of the target will be changed due to task execution. In the following discussion, we will firstly address these techniques, and then provide a detailed introduction to the PCCAIP.
@@ II-A Perception
The integration of space-based, air-based, ground-based, and sea-based networks results in a significant increase in business volume. Multiple types of sensors deployed in these networks enable real-time perception of various targets. Meanwhile, since the targets may belong to different physical processes in the real world, a substantial amount of multi-modal data are collected, which poses a great challenge in supporting services and applications with extremely low latency requirements. One feasible solution is to reduce the volume of data to decrease the latency in subsequent communication and processing stages. This is why we need to perceive the significance, meaning, and effectiveness-related information of the source data in the SAGSIN. Capturing significance can be achieved through significance-oriented sampling, while capturing the other two can be accomplished through intelligent semantic extraction, which is typically facilitated by DL, knowledge graphs (KGs), and so on. The significance-oriented sampling and semantic extraction are collectively referred to as perception in this survey.
Perception plays a significant role in various application scenarios. For example, in the emergency response and rescue applications for maritime incidents, ocean buoys deployed in the maritime area are used to collect oceanographic source data such as sea currents, wind speeds, and wave conditions. UAVs deployed above the sea surface capture image data with cameras. During non-urgent situations, the sensors operate at a relatively low sampling frequency to conserve energy and reduce data processing burden. However, in the emergency events such as maritime shipwrecks, oil spills, fires, or drowning, the sensors immediately increase their sampling frequency to keep data freshness and let the server estimate the source status more accurately. The source data are further processed to extract semantic information, such as the situation of stranded crew members and ocean current information. The extracted semantic information is then transmitted to the emergency rescue center in real-time to facilitate rescue operations.
@@ II-B Computing
In the semantic-empowered SAGSIN, computing refers to the AI-related complex operations, such as the feature extraction, semantic coding, and adaptive decision. Therefore, it is imperative to use AI algorithms to extract semantic information and make intelligent decisions, which will release data traffic and thus decrease processing latency. However, AI algorithms rely heavily on sufficient computing resources, and it is unlikely to achieve satisfactory performance when these algorithms are deployed on end devices such as ocean buoys and UAVs due to limitations in computing and caching resources. The SAGSIN encompasses various types of devices, resulting in a large volume of data that usually require real-time processing and analysis. Edge computing technology has emerged to address this issue.
In recent years, there have been significant academic and industrial interests in on-orbit edge computing based on space-based networks[40, 41], as well as edge storage technologies [27, 42]. A large number of cloud servers, edge servers, and other computing units are distributed in the SAGSIN, providing supports for perception, communication, and actuation techniques. For instance, in the context of satellite-based edge computing, efficient interconnection is achieved among space-based, and ground-based networks by leveraging LEO satellites as the core. An edge computing platform is built on the satellites, empowering the entire network with on-orbit capabilities for intelligent data collection, processing, and caching[43]. The authors assume a geographically dispersed remote team with diverse members. Through augmented reality (AR) and virtual reality (VR) technologies, remote team members can share a virtual workspace, enabling collaborative working and real-time communications. In this scenario, the on-orbit edge servers are equipped with a semantic encoding network driven by DL models. Team members wear head-mounted displays integrated with semantic decoding networks, accessing the shared virtual workspace through VR technology. In this example, the satellite-based edge computing platform provides robust support for the combination of VR technology and semantic communications.
@@ II-C Actuation
The receiver utilizes the received messages, which are relevant to task performance, to accomplish the task decision. Based on the decision results, the task execution process will exert an influence on the physical world. In this survey, the task decision and task execution comprise the definition of actuation. Taking remote control services as example, the intelligent decision-maker generates and transmits task decision commands based on the received messages, and then the terminal receives the decision commands and executes specific tasks.
With the development of 5G/6G, the variety of task-critical applications in the SAGSIN continues to enrich. For instance, the stability control of UAVs can be modeled as a classical CartPole problem solution. In this scenario, sensors collects environmental data, which will be then processed on the edge server using DL-based autoencoders to extract task-related semantic features. These features are transmitted to the remote control center on the ground to generate control commands, enabling remote control of the UAVs [44].
Another example is found in remote surgeries, which is a typical use case of haptic Internet. Sensors located in remote areas (e.g., mountainous areas and remote seas) collect crucial patient data, which are transmitted to the medical center with the aid of satellites. Surgeons receive real-time auditory, visual, and haptic feedback through a human-machine interface. Visual feedback is provided using streaming technologies, such as holographic-type communications, depending on whether surgeons interact with holograms or wear head-mounted devices. Subsequently, surgeons operate haptic devices through the human-machine interface and perform surgical actions based on real-time visual feedback and haptic information transmitted to the robot[45].
@@ II-D PCCAIP
According to the interpretation of perception, computation, actuation, we introduce the perception-communication-computing-actuation-integrated framework under the guidance of PCCAIP. This framework aims to maximize the effectiveness of tasks by establishing closed information flow loop, as shown in Fig. 3. We consider an emergency detection and rescue scenario in the SAGSIN as an example. It is worth noting that, in the current industrial infrastructure, the emergency detection task and emergency rescue task can both be independently implemented in real world. However, the collaborative design and coordination of detection & rescue operations still rely on the comprehensive SAGSIN which provides heterogeneous connections as well as computing resources.
Specifically, various sensors (e.g., cameras deployed on the UAVs) continuously observe the environment that contains abnormal statuses (e.g., sinking ships or drowning persons), and decide to sample data representing the current statuses when the data are significant enough (e.g. untimeliness, status change, or environment change). After data sampling, these data will be further processed by the computing-intensive semantic extractor deployed on UAVs and then transmitted to remote edge servers (e.g. servers deployed on maritime base station or space base station), by which multi-level data perceptions are completed by distributed computing resources.
Based on the received status data, the edge servers conduct real-time decision making and generate commands, which can guide the terminals (e.g., rescue ships) in responding to the emergencies if necessary. Subsequently, the decision commands are transmitted to remote terrestrial cloud server via LEO/GEO satellite relays. After the cloud server receives the decision commands, it will undertake task execution (e.g., rescuing the ships/persons, or doing nothing) based on the commands, by which task actuation is completed. Moreover, in cases where rescue capabilities are constrained, the cloud servers will firstly aggregate the received commands from all the edge servers and secondly decide which emergencies should be rescued according to various factors (e.g., significance, environment statuses, and effectiveness aspects). After the decision making process, the task execution will be completed by the terminal.
Task execution will further determine the statuses and thus the decisions at the next time interval (for example, if a sinking ship gets rescued by the rescue ship, there will be no emergency in the next period of time, and thus the rescue ship does not need to rescue in the next), by which closed loops for the emergency rescue tasks are formed by the PCCAIP framework.
@@ II-E Lessons Learned from This Section
This section elaborates on how the perception, computing, and actuation techniques help realize semantic communications. Overall, perception techniques help select the data with more semantics, computing techniques facilitate intelligent communications in various aspects, and actuation techniques aid in the interaction between system and environment to complete all-inclusive tasks. PCCAIP integrates all these techniques which can realize a closed-loop communication and a joint design to achieve high effectiveness of tasks. Moreover, maritime rescue tasks are adopted as an example to showcase how PCCAIP guides semantic communication system design.
The proposed PCCAIP is essentially a novel paradigm which significantly deviates from traditional integrated sensing and communication (ISAC) paradigm[46] along with integrated sensing, communication, and computing (ISCC) paradigm[47] in two aspects. Firstly, the ISAC and ISCC both focus on the integration of radar sensing (as well as traditional data capturing) and communication (computing), which belong to the category of traditional bit-level communication, while PCCAIP aims at high-level semantic communication facilitated by intelligent significance-oriented sampling and meaning extraction. Secondly, ISAC and ISCC both address the issue of only communication processes, while PCCAIP can facilitate more comprehensive task-critical applications by jointly designing both communication processes and the ultimate actuation processes. Therefore, it is a promising idea to implement semantic-communication-based task-critical applications in the future network via PCCAIP.
As typical semantic communication frameworks facilitated by PCCAIP, we will introduce significance-oriented, meaning-oriented, and effectiveness/task-oriented communications systems respectively in the following sections. In Section III, we design a significance-oriented semantic communication system which belongs to a perception-communication-integrated semantic communications. In Section IV, we study a meaning-oriented semantic communication system which falls under the category of a perception-communication-computing-integrated semantic communication system. In Section V, we introduce a task-oriented communication system which achieves perception-communication-computing-actuation integration.
@@
In this section, we will review the research on semantic communications from the data significance perspective. Here, the term “semantics” is explained as its etymological meaning “significance”, which implies a significance-oriented semantic communication system.
In significance-oriented semantic communications in the SAGSIN, a joint perception-communication framework is adopted as shown in Fig. 4. In Fig. 4, the source data can be of various types, such as given data and status updates. Given data are usually collected from fixed files provided by users, such as texts, images, and audios, while status updates are collected from a time-variant physical process from the real world. Aiming at studying the significance evolution and optimization of a physical process, all of the works reviewed in this section assume that the source data are status updates. Specifically, source data are firstly sampled by significance-oriented sampler in order to release the heave data burden, and the sampled data will be coded and modulated for transmission over wireless channel. The receiver demodulates and decodes the received message to recover sampled data. According to the recovery results, the significance of source data varies to affect the evolution of significance metric; sequentially, the variation of significance metric will affect the further decision of sampler and coder/modulator at the transmitter via feedback from the receiver. It is worth noting that both coding and modulation techniques mentioned in this section are traditional technologies based on Shannon information theory.
Considering the fundamental limits of the SAGSIN and the time-sensitivity of various tasks facilitated by SAGSIN, the data freshness/timeliness is naturally significant for semantic communication system due to large communication distance. Therefore, timeliness metrics, including age of information (AoI) and its variants, are of vital importance in describing the performance of the tasks. Thus, we will first introduce AoI and AoI variants as significance metrics in subsection III-A, and then review the literature on techniques towards significance-protection in subsection III-B.